OpenStudy (anonymous):

Product (3a+4) (2a-1)

OpenStudy (anonymous):

3a/4 + 1 = (2a - 1)/3 for now, work on the right side: divide through by 3 to make it "(2a/3) - (1/3)" this brings your equation to: 3a/4 + 1 = 2a/3 - 1/3 now you can isolate all the terms that include an a to one side, by subtracting the 2a/3 from both sides (remember what you do to one side, you must do to the other) 3a/4 - 2a/3 + 1 = 2a/3 - 2a/3 - 1/3 and that equals; 3a/4 - 2a/3 + 1 = -1/3 then subtract the 1 from both sides, in order to move it to the right side: 3a/4 - 2a/3 +1 -1 = -1/3 - 1, which equals: 3a/4 - 2a/3 = - 4/3 now use a common denominator of 12 on the left side and subtract: 3*3a/12 - 4*2a/12 = -4/3 which equals: 9a/12 - 8a/12 = - 4/3 now subtract the fractions: 1a/12 = -4/3 then multiply both sides by 12: a = 12 * (-4/3) = -16 now check the answer by plugging it back in: 3*(-16)/4 + 1 = (2*(-16) - 1)/3 gives: -11 = -11

OpenStudy (anonymous):

(3a+4) (2a-1) = 6a^2 -3a +8a -4=6a^2 +5a -4

OpenStudy (anonymous):

so finally (3a+4) (2a-1) = 6a^2 -3a +8a -4=6a^2 +5a -4